# Peter Reinhard Hansen§    and   Chen Tong

### v1

@@ B.1 MGF for ARG model
$\mathbb{E}_{T}\left(e^{zx_{T+H}}\right)$ $=\mathbb{E}_{T}\left[\mathbb{E}_{T+1}\left(e^{zX_{T+H}}\right)\right]$
$=\mathbb{E}_{T}\left[\exp\left(A(H-1,z)+\sum_{j=1}^{p}B_{i}\left(H-1,z\right)X_{T+2-j}\right)\right]$
$=\exp\left(A(h,z)+\sum_{i=2}^{p}B_{i}\left(h,z\right)x_{t+2-i}\right)\mathbb{E}_{t}\left[\exp\left(B_{1}\left(h,z\right)x_{t+1}\right)\right]$
$=\exp\left(A(h,z)+\sum_{i=1}^{p-1}B_{i+1}\left(h,z\right)x_{t+1-i}\right)$
×$\quad\times\exp\left(\frac{\eta B_{1}\left(h,z\right)}{1-\eta B_{1}\left(h,z\right)}\sum_{i=1}^{p}\beta_{i}x_{t+1-i}-\delta\log(1-\eta B_{1}\left(h,z\right))\right)$
$=\exp\left(A(h+1,z)+\sum_{i=1}^{p}B_{i}\left(h+1,z\right)x_{t+1-i}\right),$
@@
$B_{i}\left(h+1,z\right)=\begin{cases}B_{i+1}\left(h,z\right)+\frac{\eta B_{1}\left(h,z\right)}{1-\eta B_{1}\left(h,z\right)}\beta_{i}&1\leq i<p\\ \frac{\eta B_{1}\left(h,z\right)}{1-\eta B_{1}\left(h,z\right)}\beta_{i}&i=p\end{cases}$
@@ B.2 The MGF in Heston-Nandi GARCH Model
Some Computer Code
Moment of NIG distribution

### v2

@@ B.1 MGF for ARG model
$\mathbb{E}_{T}\left(e^{zx_{T+h+1}}\right)$ $=\mathbb{E}_{T}\left[\mathbb{E}_{T+1}\left(e^{zX_{T+h+1}}\right)\right]$
$=\mathbb{E}_{T}\left[\exp\left(A(h,z)+\sum_{j=1}^{p}B_{j}(h,z)X_{T+2-j}\right)\right]$
$=\exp\left(A(h,z)+\sum_{j=2}^{p}B_{j}(h,z)X_{T+2-j}\right)\mathbb{E}_{T}\left[\exp\left(B_{1}(h,z)X_{T+1}\right)\right]$
$=\exp\left(A(h,z)+\sum_{j=1}^{p-1}B_{j+1}(h,z)X_{T+1-j}\right)$
×$\quad\times\exp\left(-\delta\log\left(1-\eta B_{1}(h,z)\right)+\frac{\eta B_{1}(h,z)}{1-\eta B_{1}(h,z)}\sum_{j=1}^{p}\beta_{j}X_{T+1-j}\right)$
$=\exp\left(A(h+1,z)+\sum_{j=1}^{p}B_{j}(h+1,z)x_{T+1-j}\right),$
@@
$B_{j}\left(h+1,z\right)=\begin{cases}B_{j+1}(h,z)+\frac{\eta B_{1}(h,z)}{1-\eta B_{1}(h,z)}\beta_{j}&1\leq j<p\\ \frac{\eta B_{1}(h,z)}{1-\eta B_{1}(h,z)}\beta_{j}&j=p\end{cases}$
@@ B.2 The MGF in Heston-Nandi GARCH Model
An Direct Proof of Lemma A.1.
Proof of Lemma A.1. By the substitution, $y=zx\in\mathbb{C}$, we have
$\lim_{K\rightarrow\infty}\int_{0}^{K}x^{r}e^{-zx}\mathrm{d}x=\frac{1}{z^{r+1}}\int_{0}^{sK+itK}y^{r}e^{-y}\mathrm{d}y.$
The function $f(y)\equiv y^{r}e^{-y}$, has a pole at $y=0$ if $r<0$, and is a multi-valued function when $r$ is a non-integer. We will adopt the principal value by restricting the angle of $y=\lambda e^{i\theta}$ to $\theta\in(-\pi,\pi]$, and construct a closed contour integral that avoids the pole at $y=0$. The contour is illustrated in Figure S.1.
Figure S.1: The closed contour for the integral, $\oint y^{r}e^{-y}\mathrm{d}y$.
The function, $f(y)$, is analytic over the closed contour for $\lambda>0$, (i.e. it does not include any singularity). Therefore, by the Cauchy Integral Theorem, we have
The integral we seek is therefore given by,
This completes the proof. $\square$
@@
Code for Implementation